# Practical solutions to the relative pose of three calibrated cameras

### v4

@@ 8 Experiments on real data
In this section, we aim to further study the performance of the proposed methods, supplementing Sec. 4 (paragraph “Experiments on real data") of main paper with more detailed evaluations. Section 8.1 presents results on individual scenes, extending the analysis in Fig. 4 of the main paper. Section 8.2 investigates the impact of varying the RANSAC epipolar threshold on solver performance. These experiments extend Fig. 13 (bottom) by comparing additional solvers across all three datasets. Section 8.3 evaluates and compares the run-times of each of the proposed and state-of-the-art solvers. Section 8.4 explores the robustness of the solvers under varying inlier ratios using semi-synthetic data. In Section 8.5, we provide results using Poselib RANSAC [28] for an alternative pose error metric that considers errors across all three camera pairs. Section 8.7 evaluates the solvers within the GC-RANSAC [2] framework for all three datasets.
@@ 8.1 Results on individual scenes
Fig. 4 in the main paper showed results jointly on all PhotoTourism [23] scenes (except St. Peter’s Square), jointly on the 5 Cambridge Landmarks [24] scenes (except the Street scene, which is commonly not used due to issues with its ground truth), and Aachen Day-Night v1.1 [60]. It also showed results on one individual scene from [24], i.e., the St. Mary’s Church scene. In Fig. 15, we provide results for the accuracy-speed trade-off evaluation for all remaining individual scenes of PhotoTourism [23] and Cambridge Landmarks [24].
As discussed in the main paper, the accuracy of the proposed approximate solvers is scene-dependent. This also applies to the state-of-the-art 4p3v(HC) solver [22], since in this solver the scene needs to be similar enough to the training scenes for the MLP-based classifier to work well. The proposed ENM refitting suppresses to some extent the scene dependency of the proposed 4p3v(M)-based and 4p3v(A)-based solvers. It can be seen that the proposed 4p3v(M)+R+F solver consistently provides the best speed-accuracy trade-off both with and without ENM accross all scenes. 4p3v(A)+R+F+ENM provides a similar performance, typically beating 4p3v(HC) [22]. However, it may perform worse for some specific scenes, e.g., Shop Facade and Palace of Westminster. In general, the results on individual scenes are consistent with the results from the main paper.
@@ 8.3 Solver run-times
In this section, we present run-times of the proposed solvers as well as the state-of-the-art solvers for the relative pose problem of three calibrated cameras. While the main paper reports run-time results for full RANSAC-based estimation, we now report the run-times of the individual solvers outside of RANSAC. To measure the run-times of the solvers888Note that for 4p3v(HC) solver, in Tab 4, the time needed to load the weights of the network (or any other required data) is not added to the runtime of the solver. The data are loaded once per RANSAC, and the loading takes on average 45ms., we calculated the average run-time of each solver on more than 50k instances of the Sacre Coeur scene of the PhotoTourism dataset [52]. The run-times are reported in Table 4. The experiments were performed on an Intel(R) Core(TM) i9-10900X CPU @ 3.70GHz. In general, the implementations of all proposed solvers are not optimized for speed, and we still see room for speeding them up.
@@ 8.4 Outlier experiments
To show how the different solvers perform even under varying inlier ratios, we perform a semi-synthetic experiment. We use the Notre Dame scene from PhotoTourism [23]. We keep all inlier triplets w.r.t. a 5px epipolar threshold using the ground truth poses. We add additional synthetic outlier correspondences by generating random points in all three views. This allows us to study how the different methods perform when the inlier ratio changes. The results are shown in Fig. 17. As expected, the performance of all solvers decreases with lower inlier ratios. We also observe that 4p3v(M)+R+F performs well even with a low inlier ratio. In contrast, the relative performance of 4p3v(A)+R+F+ENM worsens with a decreased inlier ratio. However, we note that even with an inlier ratio of 40%, it still results in performance comparable to the baseline 5pt+P3P solver. This suggests that a high inlier ratio is not necessary for the ENM to work well in conjunction with the solver 4p3v(A).
| Phototourism [23] | Cambridge Landmarks [24] | Aachen Day-Night v1.1 [60] | ||||||||||||||||
@@ 8.5 Alternative evaluation measure
8.6 Comparison with [37]
The authors of [37] kindly shared their source code with us. Unfortunately, we were not able to run the part of the code that samples epipole candidates from a 10-degree polynomial curve (appropriately sampling the curve is hard as the epipole can be arbitrarily far from the image center). At the same time, the authors were also not able to run it.
Based on the working parts of the code, we tested an oracle version of [37], where instead of sampling the 10-degree polynomial curve, the oracle gives us the correct epipole. Given a sample close to the correct epipole, [37] performs comparable to our M-based solvers. In practice it is hard to find good samples (the epipole can be arbitrarily far from the image center). [37] report using 40-1,000 samples with additional local optimization for robust estimation. Even then, [37] show that this approach performs worse than 5pt+P3P on synthetic data. In contrast, two additional samples in our -based solvers already lead to better accuracy than 5pt+P3P.
| Phototourism [23] | ||||||
@@ 8.6 Comparison with [37]
| Cambridge Landmarks [24] | ||||||
| 4p3v(HC) [22] | 8.13 | 3.05 | 43.75 | 60.73 | 73.93 | 2.37 |
@@
| Aachen Day-Night v1.1 [60] | ||||||
| 4p3v(HC) [22] | 10.73 | 3.84 | 39.94 | 53.01 | 64.77 | 1.90 |
@@ 8.7 GC-RANSAC
Besides PoseLib’s RANSAC implementation, we also evaluated and compared our proposed solvers with the state-of-the-art solvers inside the GC-RANSAC [2] framework.
In GC-RANSAC, local optimization (LO) is performed using non-minimal solvers that fit models to larger-than-minimal samples. We use the non-minimal version999The non-minimal version of the 5pt solver [36] uses the last four vectors from the SVD/QR decomposition of a matrix instead of the 4-dim null space of a matrix to parameterize the unknown essential matrix. of the 5pt solver [36] and the non-minimal absolute pose DLSPnP [19] solver.101010This may not be the most efficient way how to perform non-minimal refitting. However, since all methods use the same LO, it is sufficient for a fair comparison. In contrast to LO used in Poselib RANSAC, where the estimated model is used as an initialization of the Levenberg–Marquardt algorithm, in GC-RANSAC, the estimated model is used only to score inliers.
Tab. 6 shows the results for GC-RANSAC with PROSAC sampling [6] and a 5px epipolar threshold for all scenes from the PhotoTourism [23] dataset, 5 scenes from the Cambridge Landmarks [24] dataset, and the Aachen Day-Night v1.1 [60] dataset. Similarly to what was observed for Poselib RANSAC (see Table 2 in the main paper), with the suggested modifications, all the proposed solvers outperform the state-of-the-art 4p3v(HC) solver [22] and the baseline 5pt+P3P solver in terms of pose accuracy with comparable runtimes. Again, the -based solvers provide, in general, the best speed-accuracy trade-off.
@@
Another difference is in refitting using ENM. For GC-RANSAC, the effect of ENM is not as significant as for Poselib RANSAC. When applied without refinement (+R), the early non-minimal refitting (ENM), in general, increases the precision of solvers. When combined with +R, the improvement is not very visible. The reason is that the model returned after refining the initial, approximate model estimated by the 4p3v(M)-based and 4p3v(A)-based solvers on the correspondence is usually sufficiently accurate to score inliers. Moreover, in the LO of GC-RANSAC, this approximate model is refitted using the non-minimal 5pt solver (which is similar to the refitting that is used in ENM), and the non-minimal DLSPnP solver [19]. Similarly to ENM, the filtering (+F) that uses the correspondence does not bring an improvement that is as visible as for Poselib RANSAC. This is because for GC-RANSAC, the speedup obtained using the filtering +F is not as significant, compared to the longer running times of the LO part of GC-RANSAC.
On the other hand, the remaining two suggested modifications, i.e., the -based solvers and the refinement (+R) using the correspondence bring visible improvements. This behavior is also visible in Figure 19. Here we present an ablation study on the effects of the various modifications ( and +F/+R/+ENM, which were introduced in Sec. 3.2 of the main paper) on the 4p3v(M)-based and 4p3v(A)-based solvers. The results are reported on the St. Mary’s Church, Shop Facade, and King’s College scenes from the Cambridge Landmarks dataset [24]. These results especially highlight the importance of the refinement using the point in the third view (+R). On the other hand, the benefits of 4p3v(M)-based solvers over the 4p3v(M)-based solvers that are also visible in Table 6 are not so significant as in Poselib RANSAC. Still, 4p3v(M)-based solvers lead to improved pose accuracy.

### v5

@@ 8 Experiments on real data
In this section, we aim to further study the performance of the proposed methods, supplementing Sec. 4 (paragraph “Experiments on real data”) of main paper with more detailed evaluations. Section 8.1 presents results on individual scenes, extending the analysis in Fig. 4 of the main paper. Section 8.2 investigates the impact of varying the RANSAC epipolar threshold on solver performance. These experiments extend Fig. 13 (bottom) by comparing additional solvers across all three datasets. Section 8.3 evaluates and compares the run-times of each of the proposed and state-of-the-art solvers. Section 8.4 explores the robustness of the solvers under varying inlier ratios using semi-synthetic data. In Section 8.5, we provide results using Poselib RANSAC [PoseLib] for an alternative pose error metric that considers errors across all three camera pairs. Section 8.7 evaluates the solvers within the GC-RANSAC [barath2017graph] framework for all three datasets.
@@ 8.1 Results on individual scenes
Fig. 4 in the main paper showed results jointly on all PhotoTourism [IMC2020] scenes (except St. Peter’s Square), jointly on the 5 Cambridge Landmarks [kendall2015cambridge] scenes (except the Street scene, which is commonly not used due to issues with its ground truth), and Aachen Day-Night v1.1 [zhang2021aachen]. It also showed results on one individual scene from [kendall2015cambridge], i.e., the St. Mary’s Church scene. In Fig. 15, we provide results for the accuracy-speed trade-off evaluation for all remaining individual scenes of PhotoTourism [IMC2020] and Cambridge Landmarks [kendall2015cambridge].
As discussed in the main paper, the accuracy of the proposed approximate solvers is scene-dependent. This also applies to the state-of-the-art 4p3v(HC) solver [Hruby_cvpr2022], since in this solver the scene needs to be similar enough to the training scenes for the MLP-based classifier to work well. The proposed ENM refitting suppresses to some extent the scene dependency of the proposed 4p3v(M)-based and 4p3v(A)-based solvers. It can be seen that the proposed 4p3v(M)+R+F solver consistently provides the best speed-accuracy trade-off both with and without ENM accross all scenes. 4p3v(A)+R+F+ENM provides a similar performance, typically beating 4p3v(HC) [Hruby_cvpr2022]. However, it may perform worse for some specific scenes, e.g., Shop Facade and Palace of Westminster. In general, the results on individual scenes are consistent with the results from the main paper.
| Phototourism [IMC2020] | ![]() |
| Cambridge [kendall2015cambridge] | ![]() |
| Aachen [zhang2021aachen] | ![]() |
@@ 8.3 Solver run-times
In this section, we present run-times of the proposed solvers as well as the state-of-the-art solvers for the relative pose problem of three calibrated cameras. While the main paper reports run-time results for full RANSAC-based estimation, we now report the run-times of the individual solvers outside of RANSAC. To measure the run-times of the solvers888Note that for 4p3v(HC) solver, in Tab 5, the time needed to load the weights of the network (or any other required data) is not added to the runtime of the solver. The data are loaded once per RANSAC, and the loading takes on average 45ms., we calculated the average run-time of each solver on more than 50k instances of the Sacre Coeur scene of the PhotoTourism dataset [snavely2006photo]. The run-times are reported in Table 5. The experiments were performed on an Intel(R) Core(TM) i9-10900X CPU @ 3.70GHz. In general, the implementations of all proposed solvers are not optimized for speed, and we still see room for speeding them up.
@@ 8.4 Outlier experiments
To show how the different solvers perform even under varying inlier ratios, we perform a semi-synthetic experiment. We use the Notre Dame scene from PhotoTourism [IMC2020]. We keep all inlier triplets w.r.t. a 5px epipolar threshold using the ground truth poses. We add additional synthetic outlier correspondences by generating random points in all three views. This allows us to study how the different methods perform when the inlier ratio changes. The results are shown in Fig. 17. As expected, the performance of all solvers decreases with lower inlier ratios. We also observe that 4p3v(M)+R+F performs well even with a low inlier ratio. In contrast, the relative performance of 4p3v(A)+R+F+ENM worsens with a decreased inlier ratio. However, we note that even with an inlier ratio of 40%, it still results in performance comparable to the baseline 5pt+P3P solver. This suggests that a high inlier ratio is not necessary for the ENM to work well in conjunction with the solver 4p3v(A).
| Phototourism [IMC2020] | Cambridge Landmarks [kendall2015cambridge] | Aachen Day-Night v1.1 [zhang2021aachen] | ||||||||||||||||
@@ 8.5 Alternative evaluation measure
8.6 Comparison with [DBLP:journals/ijcv/NisterS06]
The authors of [DBLP:journals/ijcv/NisterS06] kindly shared their source code with us. Unfortunately, we were not able to run the part of the code that samples epipole candidates from a 10-degree polynomial curve (appropriately sampling the curve is hard as the epipole can be arbitrarily far from the image center). At the same time, the authors were also not able to run it.
Based on the working parts of the code, we tested an oracle version of [DBLP:journals/ijcv/NisterS06], where instead of sampling the 10-degree polynomial curve, the oracle gives us the correct epipole. Given a sample close to the correct epipole, [DBLP:journals/ijcv/NisterS06] performs comparable to our M-based solvers. In practice it is hard to find good samples (the epipole can be arbitrarily far from the image center). [DBLP:journals/ijcv/NisterS06] report using 40-1,000 samples with additional local optimization for robust estimation. Even then, [DBLP:journals/ijcv/NisterS06] show that this approach performs worse than 5pt+P3P on synthetic data. In contrast, two additional samples in our -based solvers already lead to better accuracy than 5pt+P3P.
| Phototourism [IMC2020] | ||||||
@@ 8.6 Comparison with [DBLP:journals/ijcv/NisterS06]
| Cambridge Landmarks [kendall2015cambridge] | ||||||
| 4p3v(HC) [Hruby_cvpr2022] | 8.13 | 3.05 | 43.75 | 60.73 | 73.93 | 2.37 |
@@
| Aachen Day-Night v1.1 [zhang2021aachen] | ||||||
| 4p3v(HC) [Hruby_cvpr2022] | 10.73 | 3.84 | 39.94 | 53.01 | 64.77 | 1.90 |
@@ 8.7 GC-RANSAC
Besides PoseLib’s RANSAC implementation, we also evaluated and compared our proposed solvers with the state-of-the-art solvers inside the GC-RANSAC [barath2017graph] framework.
In GC-RANSAC, local optimization (LO) is performed using non-minimal solvers that fit models to larger-than-minimal samples. We use the non-minimal version999The non-minimal version of the 5pt solver [Nister-5pt-PAMI-2004] uses the last four vectors from the SVD/QR decomposition of a matrix instead of the 4-dim null space of a matrix to parameterize the unknown essential matrix. of the 5pt solver [Nister-5pt-PAMI-2004] and the non-minimal absolute pose DLSPnP [dlspnp] solver.101010This may not be the most efficient way how to perform non-minimal refitting. However, since all methods use the same LO, it is sufficient for a fair comparison. In contrast to LO used in Poselib RANSAC, where the estimated model is used as an initialization of the Levenberg–Marquardt algorithm, in GC-RANSAC, the estimated model is used only to score inliers.
Tab. 7 shows the results for GC-RANSAC with PROSAC sampling [prosac2005chum] and a 5px epipolar threshold for all scenes from the PhotoTourism [IMC2020] dataset, 5 scenes from the Cambridge Landmarks [kendall2015cambridge] dataset, and the Aachen Day-Night v1.1 [zhang2021aachen] dataset. Similarly to what was observed for Poselib RANSAC (see Table 2 in the main paper), with the suggested modifications, all the proposed solvers outperform the state-of-the-art 4p3v(HC) solver [Hruby_cvpr2022] and the baseline 5pt+P3P solver in terms of pose accuracy with comparable runtimes. Again, the -based solvers provide, in general, the best speed-accuracy trade-off.
@@
Another difference is in refitting using ENM. For GC-RANSAC, the effect of ENM is not as significant as for Poselib RANSAC. When applied without refinement (+R), the early non-minimal refitting (ENM), in general, increases the precision of solvers. When combined with +R, the improvement is not very visible. The reason is that the model returned after refining the initial, approximate model estimated by the 4p3v(M)-based and 4p3v(A)-based solvers on the correspondence is usually sufficiently accurate to score inliers. Moreover, in the LO of GC-RANSAC, this approximate model is refitted using the non-minimal 5pt solver (which is similar to the refitting that is used in ENM), and the non-minimal DLSPnP solver [dlspnp]. Similarly to ENM, the filtering (+F) that uses the correspondence does not bring an improvement that is as visible as for Poselib RANSAC. This is because for GC-RANSAC, the speedup obtained using the filtering +F is not as significant, compared to the longer running times of the LO part of GC-RANSAC.
On the other hand, the remaining two suggested modifications, i.e., the -based solvers and the refinement (+R) using the correspondence bring visible improvements. This behavior is also visible in Figure 19. Here we present an ablation study on the effects of the various modifications ( and +F/+R/+ENM, which were introduced in Sec. 3.2 of the main paper) on the 4p3v(M)-based and 4p3v(A)-based solvers. The results are reported on the St. Mary’s Church, Shop Facade, and King’s College scenes from the Cambridge Landmarks dataset [kendall2015cambridge]. These results especially highlight the importance of the refinement using the point in the third view (+R). On the other hand, the benefits of 4p3v(M)-based solvers over the 4p3v(M)-based solvers that are also visible in Table 7 are not so significant as in Poselib RANSAC. Still, 4p3v(M)-based solvers lead to improved pose accuracy.